OpenStudy (anonymous):

Differentiate the function y=(x^9-3x^4+8)/(x)

OpenStudy (dumbcow):

quotient rule \[(f/g)' = \frac{f'g -fg'}{g^{2}}\]

OpenStudy (dumbcow):

or split up fraction... \[y = x^{8}-3x^{3}+\frac{8}{x}\] then take derivative of each term

OpenStudy (anonymous):

okay thanks